# Characteristic initial value problems for the Einstein-Maxwell-scalar field equations in spherical symmetry

### v2

@@ V.1 Charged black hole solutions in null coordinates
For finding charged black hole solutions we work with the double null foliation of Sec. III.3 where the null boundary has nontrivial Misner-Sharp mass.
@@
$\begin{split}&r(0,0)=r_{0}>0\;\;,\;\;r_{,w}(0,0)=r_{N}\;\;,\;\;\\ &Q(0,0)=Q_{0}\;\;,\;\;r_{,\lambda}(0,0)=\Theta\end{split}$ (115)
@@
$Q(w,0)=$ $Q_{0}\;\;,\;\;$ (117b)
inserting these values into (116c) and integration yields
@@
Now the hypersurface equation for the given initial boundary value data are
@@
The integration of (120) with the boundary data (117), (118) and (119) gives us
@@
Consequently the most general form of metric a function of the initial values reads
@@
$ds^{2}=(r_{0}+r_{N}w)^{2}q_{AB}dx^{A}dx^{B}$ (123)
and for $w=0=dw$ that is on $\mathscr{N}^{\epsilon}(0)$
$ds^{2}=(r_{0}+\Theta\lambda)^{2}q_{AB}dx^{A}dx^{B}$ (124)
@@
$M=$ $\frac{1}{2}\left[r_{\mathcal{B}}-2\epsilon r_{\mathcal{B}}r_{N}r_{1}+Q^{2}_{0}\left(\frac{1}{r_{\mathcal{B}}}-\frac{1}{r}\right)\right]$ (125)
@@
$\lim_{w\rightarrow 0\atop\lambda\rightarrow 0}M=$ $\frac{r_{0}}{2}(1-2\epsilon r_{N}\Theta)$ (128)
@@
$\lim_{\lambda\rightarrow\infty}M=$ $\frac{r_{\mathcal{B}}}{2}\left[1-2\epsilon r_{N}r_{1}+\frac{Q^{2}_{0}}{r_{\mathcal{B}}^{2}}\right]$ (129)
@@
$r_{0,\pm}=m_{B}\pm\sqrt{m_{B}^{2}-Q^{2}_{0}}$ (131)
where the numerical value $r_{0,+}$ is the same value as the value of the radius of the outer event horion of a Reissner-Nordström black hole and $r_{0,-}$ the value of the inner event horizon.
V.1.1 Sub-extremal black holes
Suppose $r_{N}=0$, then the areal distance has the form
@@ V.1.1 Sub-extremal black holes
Since $w$ parameterises affine geodesics on the boundary $\mathcal{B}^{\epsilon}$ where $\lambda=0$, we have the freedom of choice in the origin of the affine parameter. The transformation
$w\rightarrow\tilde{w}=w-\epsilon\Theta\frac{2r_{0}^{3}}{Q_{0}^{2}-r_{0}^{2}}$ (133)
@@
$r=\tilde{r}(\lambda\tilde{w}):=r_{0}-\epsilon\frac{Q^{2}_{0}-r^{2}_{0}}{2r_{0}^{3}}(\tilde{w}\lambda)$ (134a)
@@
$\equiv\lambda^{2}\tilde{V}(\tilde{w}\lambda)$ (134b)
@@
$\begin{split}ds^{2}=-\lambda^{2}\tilde{V}(\lambda w)dw^{2}+2\epsilon dwd\lambda+\tilde{r}^{2}(\lambda w)q_{AB}dx^{A}dx^{B}\end{split}$ (135)
is invariant under the transformation $(w,\lambda)\rightarrow(aw,a^{-1}\lambda)$ for any number $a$, which implies a Killing vector of the form $\xi=\epsilon k(w\partial_{w}-\lambda\partial_{\lambda})$, where $k$ is a positive normalization constant determined from the calculation of the norm $\xi^{a}\xi_{a}$ and the factor $\epsilon$ is included in order to choose a future timelike Killing vector field. In accordance with the results of Sec. 95, this also gives us $A(w)=kw$ and $B=0$ for the two $w-$dependent functions of the general Killing vector in equation (95). Calculation of the asymptotic limits of the norms gives us
@@
Therefore, the Killing vector is an asymptotic timelike unit Killing vector if the Killing vector has the form
@@
As mentioned above, if the norm of the Killing vector, $\xi^{a}\xi_{a}$ vanishes for $\lambda=0$, the null hypersurface $\mathcal{B}^{\epsilon}$ is a Killing horizon. If $\epsilon=1$, this boundary is the future Killing horizon $H^{+}$, and the past event horizon $H^{-}$ for the choice $\epsilon=-1$. Its surface gravity $\kappa_{H}$ can be found from the invariant relation Wald (1984)
@@
Note also that the metric (135) is regular at the null surface $w=0$ given by $\mathcal{N}^{\epsilon}(0)$. This surface is the future (past) Killing horizon $H^{+}$ ($H^{-}$) if $\epsilon=-1$ ($\epsilon=1$) with Killing vector field $\xi=\kappa_{H}\lambda\partial_{\lambda}$ ($\xi=\kappa_{H}w\partial_{w}$). On the other hand, $\xi$ vanishes at the 2-sphere given by $w=\lambda=0$, which constitutes the bifurcating Killing horizon.
@@ V.1.2 Extremal black holes
The surface gravity $\kappa_{H}$ in (140) vanishes if $Q_{0}=m_{B}=r_{0}$, but this goes along with the vanishing of the Killing vector in (138). In fact, in that limit the resulting metric associated to the $\{w,\lambda\}$ coordinates introduced in the last subsection is not the extremal Reissner-Nordström metric, but the so called Bertotti-Robinson metric Gallo et al. (2021). Thus the Killing vector (138) cannot be used to find an extremal Reissner-Nordström black hole solution where the global charge to mass ratio is unity. Indeed, as we will show here, the resulting extremal metric is not invariant under the transformation $(w,\lambda)\rightarrow(aw,a^{-1}\lambda)$, and therefore the resulting Killing vector field can not be written as in (138).
To single out an extremal black hole solution in the general solution (121) for $r$ and $V$ which also has a non-vanishing Killing vector at the horizon of the black hole, consider (130) while setting $m_{e}:=Q_{0}=m_{B}.$ This gives
$m_{e}=\frac{m_{e}^{2}+r_{0}^{2}}{2r_{0}}-\epsilon r_{0}r_{N}\Theta$ (142)
@@
Since in the Schwarzschild case $r_{0}$ equates twice the Bondi mass, it is desirable to relate $r_{0}$ one-to-one with $m_{e}$. In (143) this is achieved only by either setting $\Theta=0$ or $r_{N}=0$ giving in both cases $r_{0}=m_{e}$. In (144), the choice $r_{0}=m_{e}$ gives $r_{N}=0$ and $\Theta\neq 0$ left arbitrary, while for (145) the same choice yields $\Theta=0$ and $r_{N}\neq 0$ left arbitrary.
Thus an extremal black hole solution can be found if $r_{0}=m_{e}=Q_{0}$ and either $r_{N}=0$ and $\Theta\neq 0$ or $\Theta=0$ and $r_{n}\neq 0$. Now consider the solution for the areal distance and metric potential $V$ for these for cases: for the case $(r_{N}=0\;,\;\Theta\neq 0)$ we find
@@
$V=$ $\left(\frac{\lambda}{m_{e}+\Theta\lambda}\right)^{2}$ (146b)
@@
The line element of first solution is
@@
but in fact it looks more familiar, if $r(\lambda)$ in (146a) is inverted to $\lambda(r)$, then we find
@@
so that the coordinate transformation $\hat{w}=w/\Theta$, $\lambda=(r-m_{e})/\Theta$ leaves us with the well known line element
@@
of an extremal Reissner-Nordström black hole with either ingoing($\epsilon=1$) or outgoing polar null coordinates. In such case, $\hat{w}$ is usually represented as $v$, and the past event horizon is reached as $v\to-\infty.$ For $\epsilon=1$, this metric covers the interior region of the past event horizon and the exterior region (in that case $\hat{w}$ is identified with the retarded null coordinate $u$). This coordinate does not cover the future horizon, which is asymptotically reached as $u\to\infty$. Indeed, calculation of the outgoing expansion rate of the surface forming null vector $\ell^{a}\partial_{a}=\partial_{\lambda}$ using (149) gives $2\Theta/(m_{e}+\Theta\lambda)$ which does not vanish for any value of $w$. Thus there is no null surface $w=const$ labelling the future event horizon. In a similar way, for ingoing null coordinates ($\epsilon=1$), the resulting coordinates cover the exterior and interior of the future event horizon.
@@
Consider now the case $\Theta=0$ which gives (147) and (147) for the non trivial components of the metric.
@@
Note also, that the surface $w=0$ is a regular null surface determined by $r=m_{e}$ and where $V=\frac{\lambda^{2}}{m^{2}_{e}}$. On the other hand, from (152) and (107), we find the following asymptotic behaviour of the Killing vector near $w=0$
@@
The asymptotic norm of the Killing vector gives a timelike unit vector by construction. Observe the difference to the nonextremal case (138), the null surface $\lambda=0$ is not a Killing horizon for (155). In fact, the boundary surface has an expansion rate proportional to $r_{N}$. An additional surface (to $w=0$) where the Killing vector becomes null is at
$\lambda(w)=-\frac{2\epsilon}{w(w+m_{e})^{2}}.$ (156)
@@
Indeed, the parameter $r_{N}$ in (147) and (148) can be eliminated using the scaling relations $w\rightarrow w/r_{N}$ and $\lambda\rightarrow r_{N}\lambda$, so that the final solution for the extremal Reissner-Nordström black hole only depends on the physical parameter $m_{e}$, the Bondi mass, so that
$r=$ $m_{e}+w+\epsilon\left[\frac{w^{2}}{2(w+m_{e})^{2}}\right]\lambda$ (157a)
$V=$ $-\frac{4m_{e}\lambda^{2}}{(m_{e}+w)}\left[\frac{\epsilon\lambda w^{3}-(m_{e}-2w)(m_{e}+w)^{3}}{\lambda^{2}w^{4}+4\epsilon\lambda w^{2}(m_{e}+w)^{3}+4(m_{e}+w)^{6}}\right]$ (157b)
@@
This solution was also obtained by us in Gallo et al. (2021) using an alternative hierarchical system of equations, note however that the $w$ used here must be identified with the $-w$ used in that reference. In difference to the traditional outgoing/ingoing Eddington-Finkelstein coordinates, the $\{w,\lambda,x^{A}\}$ is a regular coordinate system at both horizons.
@@ V.2 FJNW solution in null coordinates
The simplest Killing vector (107) is where $A=1$ and $B=0$. The Killing vector becomes $\xi=\partial_{w}$ and its additional requirement to be asymptotically timelike restricts the asymptotic frame to be a Bondi frame, i.e. $H=1$ so that $r\sim\lambda$ for large values of $\lambda$. Hereafter, we consider a real and uncharged scalar field solution possessing this symmetry. We will find the FJNW solution Fisher (1948); Janis et al. (1968); Wyman (1981); Virbhadra (1997); Ota et al. (2022) in affine null coordinates, which has not been derived explicitly as a solution of an hierarchical system of the Einstein-scalar field equations in affine null coordinates.
@@
where we applied the asymptotic condition, $\lim_{r\rightarrow\infty}\Phi=0$, for the scalar field. As it is also shown in (164), the parameter $a$ is the scalar monopole. Although $\Phi$ in (164) is a test scalar field, we can use this ansatz for the scalar field initial data to solve the complete Einstein-scalar field equations in a consistent way . In particular, note that if $r^{2}V$ has the same form for the complete system like in the Schwarzschild case, the scalar field will have the same form as in Schwarzschild. The solution of (158) gives us
@@
Reality of $\delta$ requires $2\kappa a^{2}<b^{2}$. Note that if $a=0$, the scalar field vanishes, and therefore the solution must reduce to the Schwarzschild metric. In that case (165) reduces to
@@
$r=\lambda+\frac{b}{2}\left(1-\delta\right)+O(\lambda^{-1})$ (168)
@@
Next, using the notation as in (36d), integration of (159) yields $Z(\lambda)=Z_{\infty}=const,$ which is related to the Bondi mass by (41), i.e.
@@
$\begin{split}V=&V_{1}\lambda+V_{0}+\frac{2\lambda}{b(\delta-1)}\left(1+\frac{b}{\lambda}\right)^{\delta}\\ &+\frac{2Z_{\infty}\left[b(\delta-1)-2\lambda\right]}{b^{2}(\delta^{2}-1)}\left(1+\frac{b}{\lambda}\right)^{\delta}\end{split}$ (173)
@@
$\begin{split}V=&\left[V_{1}+\frac{(2\delta+2)b-4Z_{\infty}}{b^{2}(\delta^{2}-1)}\right]\lambda+\left[V_{0}+\frac{2b\delta-2Z_{\infty}}{(\delta-1)b}\right]\\ &+O(\lambda^{-1})\end{split}$ (174)
@@
$V_{0}=$ $\frac{2Z_{\infty}-(1+\delta)b}{(\delta-1)b}$ (175)
$V_{1}=$ $\frac{2(1+\delta)b-4Z_{\infty}}{b^{2}(1-\delta^{2})}$ (176)
@@
Finally, the asymptotic expansion of (177) which gives
@@
By replacing this values into (177) we arrive to our final expression for $V$

### v3

@@ V.1 Charged black hole solutions in null coordinates
For finding charged black hole solutions, we work with the double-null foliation of Sec. III.3, where the null boundary has nontrivial Misner-Sharp mass.
@@
$\begin{split}&r(0,0)=r_{0}>0\;\;,\;\;r_{,w}(0,0)=r_{N}\;\;,\;\;\\ &Q(0,0)=Q_{0}\;\;,\;\;r_{,\lambda}(0,0)=\Theta.\end{split}$ (115)
@@
$Q(w,0)=$ $Q_{0}\;\;.\;\;$ (117b)
Insert these values into (116c) and integration yields
@@
Now the hypersurface equations for the given initial boundary value data are
@@
The integration of (120) with the boundary data (117), (118), and (119) gives us
@@
Consequently, the most general form of the metric for these initial values reads
@@
$ds^{2}=(r_{0}+r_{N}w)^{2}q_{AB}dx^{A}dx^{B},$ (123)
and on $\mathscr{N}^{\epsilon}(0)$, where $w=0=dw$,
$ds^{2}=(r_{0}+\Theta\lambda)^{2}q_{AB}dx^{A}dx^{B}.$ (124)
@@
$M=$ $\frac{1}{2}\left[r_{\mathcal{B}}-2\epsilon r_{\mathcal{B}}r_{N}r_{1}+Q^{2}_{0}\left(\frac{1}{r_{\mathcal{B}}}-\frac{1}{r}\right)\right],$ (125)
@@
$\lim_{w\rightarrow 0\atop\lambda\rightarrow 0}M=$ $\frac{r_{0}}{2}(1-2\epsilon r_{N}\Theta),$ (128)
@@
$\lim_{\lambda\rightarrow\infty}M=$ $\frac{r_{\mathcal{B}}}{2}\left[1-2\epsilon r_{N}r_{1}+\frac{Q^{2}_{0}}{r_{\mathcal{B}}^{2}}\right],$ (129)
@@
$r_{0,\pm}=m_{B}\pm\sqrt{m_{B}^{2}-Q^{2}_{0}},$ (131)
where the numerical value $r_{0,+}$ is the same value as the value of the radius of the outer event horizon of a Reissner-Nordström black hole and $r_{0,-}$ the value of the inner event horizon.
V.1.1 Subextremal black holes
Suppose $r_{N}=0$; then the areal distance has the form
@@ V.1.1 Subextremal black holes
Since $w$ parametrizes affine geodesics on the boundary $\mathcal{B}^{\epsilon}$ where $\lambda=0$, we have the freedom of choice in the origin of the affine parameter. The transformation
$w\rightarrow\tilde{w}=w-\epsilon\Theta\frac{2r_{0}^{3}}{Q_{0}^{2}-r_{0}^{2}},$ (133)
@@
$r=\tilde{r}(\lambda\tilde{w}):=r_{0}-\epsilon\frac{Q^{2}_{0}-r^{2}_{0}}{2r_{0}^{3}}(\tilde{w}\lambda),$ (134a)
@@
$\equiv\lambda^{2}\tilde{V}(\tilde{w}\lambda),$ (134b)
@@
$\begin{split}ds^{2}=-\lambda^{2}\tilde{V}(\lambda w)dw^{2}+2\epsilon dwd\lambda+\tilde{r}^{2}(\lambda w)q_{AB}dx^{A}dx^{B},\end{split}$ (135)
is invariant under the transformation $(w,\lambda)\rightarrow(aw,a^{-1}\lambda)$ for any number $a$, which implies a Killing vector of the form $\xi=\epsilon k(w\partial_{w}-\lambda\partial_{\lambda})$, where $k$ is a positive normalization constant determined from the calculation of the norm $\xi^{a}\xi_{a}$ and the factor $\epsilon$ is included in order to choose a future timelike Killing vector field. In accordance with the results of Sec. 95, this also gives us $A(w)=kw$ and $B=0$ for the two $w-$dependent functions of the general Killing vector in Eq. (95). Calculation of the asymptotic limits of the norms gives us
@@
Therefore, the Killing vector is an asymptotic timelike unit vector if the Killing vector has the form
@@
As mentioned above, if the norm of the Killing vector $\xi^{a}\xi_{a}$ vanishes for $\lambda=0$, the null hypersurface $\mathcal{B}^{\epsilon}$ is a Killing horizon. If $\epsilon=1$, this boundary is the future Killing horizon $H^{+}$, and the past event horizon $H^{-}$ is for the choice $\epsilon=-1$. Its surface gravity $\kappa_{H}$ can be found from the invariant relation Wald (1984)
@@
Note also that the metric (135) is regular at the null surface $w=0$ given by $\mathscr{N}^{\epsilon}(0)$. This surface is the future (past) Killing horizon $H^{+}$ ($H^{-}$) if $\epsilon=-1$ ($\epsilon=1$) with Killing vector field $\xi=\kappa_{H}\lambda\partial_{\lambda}$ ($\xi=\kappa_{H}w\partial_{w}$). On the other hand, $\xi$ vanishes at the two-sphere given by $w=\lambda=0$, which is a bifurcating Killing horizon.
@@ V.1.2 Extremal black holes
The surface gravity $\kappa_{H}$ in (140) vanishes, if $Q_{0}=m_{B}=r_{0}$, but this goes along with the vanishing of the Killing vector in (138). In fact, in that limit the resulting metric associated with the $\{w,\lambda\}$ coordinates introduced in the last subsection is not the extremal Reissner-Nordström metric, but the so-called Bertotti-Robinson metric Gallo et al. (2021). Thus the Killing vector (138) cannot be used to find an extremal Reissner-Nordström black hole solution where the global charge-to-mass ratio is unity. Indeed, as we will show here, the resulting extremal metric is not invariant under the transformation $(w,\lambda)\rightarrow(aw,a^{-1}\lambda)$, and therefore the resulting Killing vector field cannot be written as in (138).
To single out an extremal black hole solution in the general solution (121) for $r$ and $V$, which also has a nonvanishing Killing vector at the horizon of the black hole, consider (130), while setting $m_{e}:=Q_{0}=m_{B}.$ This gives
$m_{e}=\frac{m_{e}^{2}+r_{0}^{2}}{2r_{0}}-\epsilon r_{0}r_{N}\Theta,$ (142)
@@
Since in the Schwarzschild case $r_{0}$ equals twice the Bondi mass, it is desirable to relate $r_{0}$ one-to-one with $m_{e}$. In (143) this is achieved only by setting either $\Theta=0$ or $r_{N}=0$ giving in both cases $r_{0}=m_{e}$. In (144), the choice $r_{0}=m_{e}$ gives $r_{N}=0$ and $\Theta\neq 0$ left arbitrary, while for (145) the same choice yields $\Theta=0$ and $r_{N}\neq 0$ left arbitrary.
Thus an extremal black hole solution can be found if $r_{0}=m_{e}=Q_{0}$ and either $r_{N}=0$ and $\Theta\neq 0$ or $\Theta=0$ and $r_{N}\neq 0$. Now consider the solution for the areal distance and metric potential $V$ for these cases: for the case $(r_{N}=0\;,\;\Theta\neq 0)$ we find
@@
$V=$ $\left(\frac{\lambda}{m_{e}+\Theta\lambda}\right)^{2},$ (146b)
@@
The line element of the first solution is
@@
but, in fact, it looks more familiar if $r(\lambda)$ in (146a) is inverted to $\lambda(r)$, and then we find
@@
so that the coordinate transformation $\hat{w}=w/\Theta$, $\lambda=(r-m_{e})/\Theta$ leaves us with the well-known line element
@@
of an extremal Reissner-Nordström black hole with either ingoing ($\epsilon=1$) or outgoing ($\epsilon=-1$) polar null coordinates. In the first case, $\hat{w}$ is usually represented as $v$, and the past event horizon is reached as $v\to-\infty.$ For $\epsilon=1$, this metric covers the interior region of the past event horizon and the exterior region (in that case $\hat{w}$ is identified with the retarded null coordinate $u$). This coordinate does not cover the future horizon, which is asymptotically reached as $u\to\infty$. Indeed, calculation of the outgoing expansion rate of the surface forming null vector $\ell^{a}\partial_{a}=\partial_{\lambda}$ using (149) gives $2\Theta/(m_{e}+\Theta\lambda)$ which does not vanish for any value of $w$. Thus there is no null surface $w=$const labeling the future event horizon. In a similar way, for ingoing null coordinates ($\epsilon=1$), the resulting coordinates cover the exterior and interior of the future event horizon.
@@
Consider now the case $\Theta=0$, which gives (147) and (148) for the non trivial components of the metric.
@@
Also note the surface $w=0$ is a regular null surface determined by $r=m_{e}$ and where $V=\frac{\lambda^{2}}{m^{2}_{e}}$. From (152) and (107), we find the following asymptotic behavior of the Killing vector near $w=0$
@@
The asymptotic norm of the Killing vector gives a timelike unit vector by construction. Observe the difference to the nonextremal case (138), where the null surface $\lambda=0$ is not a Killing horizon for (155). In fact, the boundary surface has an expansion rate proportional to $r_{N}$. An additional surface (to $w=0$) where the Killing vector becomes null is at
$\lambda(w)=-\frac{2\epsilon(r_{N}w+m_{e})^{2}}{w}.$ (156)
@@
Indeed, the parameter $r_{N}$ in (147) and (148) can be eliminated using the scaling relations $w\rightarrow w/r_{N}$ and $\lambda\rightarrow r_{N}\lambda$, so that the final solution for the extremal Reissner-Nordström black hole depends only on the physical parameter $m_{e}$, the Bondi mass, so that
$r=$ $m_{e}+w+\epsilon\left[\frac{w^{2}}{2(w+m_{e})^{2}}\right]\lambda,$ (157a)
$V=$ $-\frac{4m_{e}\lambda^{2}}{(m_{e}+w)}\left[\frac{\epsilon\lambda w^{3}-(m_{e}-2w)(m_{e}+w)^{3}}{\lambda^{2}w^{4}+4\epsilon\lambda w^{2}(m_{e}+w)^{3}+4(m_{e}+w)^{6}}\right],$ (157b)
@@
This solution was also obtained by us in Gallo et al. (2021) using an alternative hierarchical system of equations; note, however, that the $w$ used here must be identified with the $-w$ used in that reference. In difference to the traditional outgoing/ingoing Eddington-Finkelstein coordinates, the $\{w,\lambda,x^{A}\}$ is a regular coordinate system at both horizons.
@@ V.2 FJNW solution in null coordinates
The simplest Killing vector (107) has $A=1$ and $B=0$. The Killing vector becomes $\xi=\partial_{w}$ and its additional requirement to be asymptotically timelike restricts the asymptotic frame to be a Bondi frame, i.e. $H=1$ so that $r\sim\lambda$ for large values of $\lambda$. Hereafter, we consider a real and uncharged scalar field solution possessing this symmetry. We will find the FJNW solution Fisher (1948); Janis et al. (1968); Wyman (1981); Virbhadra (1997); Ota et al. (2022) in affine-null coordinates, which has not been derived explicitly as a solution of a hierarchical system of the Einstein-scalar field equations in affine-null coordinates.
@@
where we applied the asymptotic condition, $\lim_{r\rightarrow\infty}\Phi=0$, for the scalar field. As it is also shown in (164), the parameter $a$ is the scalar monopole. Although $\Phi$ in (164) is a test scalar field, we can use this ansatz for the scalar field initial data to solve the complete Einstein-scalar field equations in a consistent way. In particular, note that if $r^{2}V$ has the same form for the complete system as in the Schwarzschild case, the scalar field will have the same form as in the Schwarzschild case. The solution of (158) gives us
@@
The reality of $\delta$ requires $2\kappa a^{2}<b^{2}$. Note that if $a=0$, the scalar field vanishes, and therefore the solution must reduce to the Schwarzschild metric. In that case (165) reduces to
@@
$r=\lambda+\frac{b}{2}\left(1-\delta\right)+O(\lambda^{-1}),$ (168)
@@
Next, using the notation as in (36d), the integration of (159) yields $Z(\lambda)=Z_{\infty}$=const, which is related to the Bondi mass by (41), i.e.
@@
$\begin{split}V=&V_{1}\lambda+V_{0}+\frac{2\lambda}{b(\delta-1)}\left(1+\frac{b}{\lambda}\right)^{\delta}\\ &+\frac{2Z_{\infty}\left[b(\delta-1)-2\lambda\right]}{b^{2}(\delta^{2}-1)}\left(1+\frac{b}{\lambda}\right)^{\delta}.\end{split}$ (173)
@@
$\begin{split}V=&\left[V_{1}+\frac{(2\delta+2)b-4Z_{\infty}}{b^{2}(\delta^{2}-1)}\right]\lambda+\left[V_{0}+\frac{2b\delta-2Z_{\infty}}{(\delta-1)b}\right]\\ &+O(\lambda^{-1}).\end{split}$ (174)
@@
$V_{0}=$ $\frac{2Z_{\infty}-(1+\delta)b}{(\delta-1)b},$ (175)
$V_{1}=$ $\frac{2(1+\delta)b-4Z_{\infty}}{b^{2}(1-\delta^{2})}.$ (176)
@@
Finally, from the asymptotic expansion of (177), which gives
@@
By replacing these values into (177) we arrive to our final expression for $V$